Planning week 13-20 November 2010

There were several ARs on the disk during ths week. The AR 11121+23 complex had produced M-flares the previous week, and gave some C-flares early this week but then faded. AR 11124 had good-sized sunspots but didn't produce much activity. AR 11126 appeared in the SE quadrant and was unusual in having old cycle magnetic polarity. It didn't produced much activity, though. Towards the end of the week, the old decaying region AR 11127 appeared around the limb with a sunspot and bright fan loops.

The week was divided into three OP periods: OP1 was 3 days long, and OP2 and OP3 were 2 days long.

Key observations

  1. HOP 174 - coordination with IBIS; various AR targets (13-17 Nov).
  2. HOP 130 - full disk slot scan (16 Nov).
  3. HOP 176 - jets in ARs; AR 11126 (19, 20 Nov).
  4. PRY's bright point trigger study for loop footpoints.
  5. PRY's sit-and-stare study for active regions.
  6. Interchange reconnection between AR and CH (18 Nov).
  7. Test of flare hunter study (14 Nov).
  8. Loop morphology and oscillations (19 Nov).

Saturday 13 November

AIA 193 image at 23:35 UT
GOES plot (13-16 Nov)

AR 11121 and 11123, a nearby pair of active regions that had produced significant flare activity around 5-7 Nov were just past the meridian at this time. A test of the bright point trigger was performed at 10:33 using study Hunter_loop_fe8 over AR 11123 to find bright footpoints.

HOP 174 (Madjarska & Doyle) was supported between 15:00 and 19:00. The pointing was set to the trailing sunspot of AR 11121. Note this observation used a lot of the EIS telemetry.

A test of the flare hunter study (FLARE266_TEST1) was performed beginning at 21:00 and continuing for 9 hours. The target was again AR 11121.

Sunday 14 November

AIA 193 image at 23:37 UT

The flare hunter study test was completed at 06:00. Note that a C1 flare occurred at 00:00 which have been caught with the study.

HOP 174 was supported again during 15:00-19:00, this time for AR 11124 in the northern hemisphere. This region had a bigger spot.

Monday 15 November

AIA 193 image at 23:36 UT

No EIS observations were performed this day.

Tuesday 16 November

AIA 193 image at 20:39 UT
GOES plot, 16-18 Nov

HOP 130 (full Sun slot scan) was supported during 10:45 - 14:17 UT. Note that HOP 179 was run afterwards, but EIS did not support.

Wednesday 17 November

AIA 193 image at 23:34 UT

HOP 174 was supported again during 15:00-18:00, this time the pointing was at AR 11126 in the SE quadrant.

At 20:10 began a long run of PRY_sit_stare_1 which was pointed above the limb at the loop system of the AR 11121+11123 complex.

Thursday 18 November

AIA 193 image
GOES plot 18-20 Nov

The PRY_sit_stare_1 run was continued until 05:16.

A study of interchange reconnection between a small, low latitude coronal hole and a weak active region was performed using dhb_polar_scan_Q90 between 14:00 and 18:00.

Another PRY_sit_stare_1 run started at 22:23, this time on AR 11124. The pointing was set at the footpoints of some fan loops, and is not off-limb.

Friday 19 November

AIA 193 image

The PRY_sit_stare_1 sequence continued until 05:21.

At 09:00 HOP 176 (small-scale jets on the solar disk) was supported for 3 hours, pointed at AR 11126. Note that this region has an old cycle polarity.

A loop morphology and oscillation study (de Moortel) was scheduled for 18:18-21:58, pointed at AR 11127.

Saturday 20 November

AIA 193 image

The Hunter_loop_fe8 was scheduled again at 01:55, with the aim of finding footpoints in AR 11127.

HOP 176 was supported again between 09:00 and 12:00 for AR 11126.